OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6. y = one thirdx - 8 y = -3x + 12 y = 3x - 24 y = -one thirdx - 4

OpenStudy (anonymous):

last one

OpenStudy (angela210793):

perpendicular----> m=-1/m1=-1/3 y-y1=m(x-x1) y+6=-1/3(x-6) y=-x/3 -4